OpenStudy (anonymous):

what is.45 as a fraction in simpilst form

OpenStudy (anonymous):

45 / 100 / 5 9 / 10

OpenStudy (anonymous):

45/100 = 9/20

OpenStudy (anonymous):

oops * 9 / 20

OpenStudy (antonio_xx2):

.45=45/100=45/5=100/5=9/20